OpenStudy (anonymous):

Write the answer in scientific notation: (5.7 x 10^(-6))(6.2 x 10^(8))

OpenStudy (anonymous):

\[3.534*10^3 \]

OpenStudy (lalaly):

\[\large{(5.7 \times 6.2)} \times 10^{-6+8}=35.34 \times 10^2\]